# Does Graph Prompt Work? A Data Operation Perspective with Theoretical Analysis

### v1

@@ 5.3 On mapping to $\epsilon\text{-}B_{G}$ with single graph
(a) Feature Number vs. epsilon
(b) Graph Size vs. epsilon
(c) Layer Number vs. epsilon
@@
Subsequently, we calculate the mean and standard deviation of these upper bounds to generate plots. The parameters of interest include rank loss (as shown in Figure 1 in section 4.1), node feature dimension (Figure 4a), graph size (Figure 4b), and model layer number (Figure 4c). Intuitively, with the increase of data complexity (e.g., larger features and graph size), the upper bound becomes larger in general. As the graph model becomes more complicated (e.g., layer number increase), the projected space becomes larger making the error bound intend to be smaller. When weight rank declines, the model’s capacity intends to poor results, making the error bound increase. More advanced graph prompts (e.g., All-in-One) usually have a lower error bound than the naive one (e.g., GPF). These observations can be naturally inferred from our theoretical analysis in section 4.1.
@@ 5.4 On mapping to $\epsilon\text{-}B_{G}$ with multiple graphs
Theorem 6 discusses a lower bound on the RMSE over the errors on multiple graphs with a single prompt token. In this section, we conducted experiments on the number of graphs in the dataset w.r.t the empirical minimum error. As shown in Figure 5a, the minimum error shows an upward trend and then tends to saturate, which is highly consistent with the findings in Theorem 6.
(a) Simple Prompt
(b) GPF Plus
(c) All in one
@@
Theorem 7 suggests that for the graph prompt with multiple tokens and multiple graphs, a small $k$ tokens is sufficient to achieve good performance. In particular, we wish to see how the error (loss) changes as $M$ (number of graphs) increases while $k$ remains fixed, and its counterpart case: how does the error change as $k$ increases while $M$ remains fixed? Here we explored the relationship between the number of prompt tokens, the number of graphs in the dataset, and the error. We present experimental results in Figure 5b and Figure 5c, which indicate two surfaces. From these figures we can find that both GPF and All-in-One show similar effects: when the number of prompt tokens exceeds $10$, the error becomes relatively small. As the number of prompt tokens increases further, the loss does not significantly decrease. Similarly, when the number of graphs increases and the number of prompt tokens is large, the decrease in error is also not that obvious.
@@ A.1.1 Preliminaries
Graph Convolutional Networks (GCN) Kipf & Welling (2016) perform convolution operations on graph-structured data by aggregating feature information from a node’s neighbors. The recursive update rule for a GCN layer is:
$\mathbf{H}^{(l+1)}=\sigma\left(\tilde{\mathbf{A}}\mathbf{H}^{(l)}\mathbf{W}^{(l)}\right),$ (7)
where $\tilde{\mathbf{A}}=\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ is the normalized adjacency matrix, $\mathbf{H}^{(l)}$ is the node feature matrix at layer $l$, $\mathbf{W}^{(l)}$ is the learnable weight matrix, and $\sigma$ is a nonlinear activation function.
Graph Attention Networks (GAT) Veličković et al. (2017) enhance GCNs by introducing attention mechanisms to weigh the importance of neighboring nodes during aggregation. In the simplest form of self-attention, the attention coefficient between node $i$ and node $j$ is based on the inner product of their feature vectors:
$e_{ij}$ $=\mathbf{X}_{i}^{\top}\mathbf{X}_{j},$ (8)
$\alpha_{ij}$ $=\frac{\exp(e_{ij})}{\sum_{k\in\mathcal{N}(i)}\exp(e_{ik})},$ (9)
$\mathbf{H}_{i}^{(l+1)}$ $=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{H}_{j}^{(l)}\mathbf{W}^{(l)}\right),$ (10)
where $\mathcal{N}(i)$ denotes the neighbors of node $i$, $\alpha_{ij}$ is the normalized attention coefficient, and $\mathbf{W}^{(l)}$ is the learnable weight matrix at layer $l$. This mechanism allows the model to focus on the most relevant neighbors when updating node representations.
@@
Pre-training and Fine-tuning Zhu et al. (2021)involves two stages: first, a model is pre-trained on a large dataset to learn general representations, and then it is fine-tuned on a specific downstream task.
@@
$F_{\theta^{*}\rightarrow\theta^{\#}}(G_{\text{ori}})\approx C(G_{\text{ori}}).$ (11)
@@
Graph Prompt Frameworks introduce learnable modifications to input graphs to enhance the performance of frozen pre-trained GNNs on downstream tasks. Two primary frameworks are prompt token vectors like GPF/GPF-Plus Fang et al. (2022) and prompt subgraph like All-in-One Sun et al. (2023a).
GPF (Graph Prompt Feature) adds a prompt vector to each node’s feature vectors. Let $\bm{\omega}=\{p\}$, $p\in\mathbb{R}^{F\times 1}$ be the learnable prompt vector, then the updated node features are:
@@
$p^{*}=\arg\min_{p}\sum_{G\in\mathcal{G}}\mathcal{L}_{T_{\text{dow}}}\left(F_{\theta^{*}}(P_{\bm{\omega}}(G))\right)$ (14)
GPF-Plus adds a combination of multiple prompt vectors to each node’s features. Let $\bm{\omega}=\{p_{1},\cdots,p_{k},Q\}$, $p_{i}\in\mathbb{R}^{F\times 1}$ be the learnable prompt vector, $Q\in\mathbb{R}^{M\times k}$. Let $P=\begin{pmatrix}p_{1}^{\top}\\ \vdots\\ p_{k}^{\top}\end{pmatrix}$. The node features are updated in such a way:
@@
All-in-One optimizes the prompt tokens and their connections to adapt the pre-trained model to downstream tasks without altering the model parameters.
@@
Diffusion Matrix Gasteiger et al. (2019) plays a crucial role in representing the diffusion process on a graph. Specifically, many GNN architectures can be expressed using the following formulation:
@@ A.1.2 Definitions
Table 2: Glossary
Term Explanation
@@
Adjacency matrix A square matrix used to represent a finite graph, where $A_{ij}=1$ if there is an edge from vertex $i$ to vertex $j$, and 0 otherwise
Diffusion matrix The matrix equivalent to graph aggregation in GNNs
Span For a set $V$, we say “p spans V” if p can take any value in V
$i$-th order embedding matrix The embedding matrix after $i$ iterations of message passing and aggregation in a GNN
Cone A set $C$ such that for any $x\in C$ and $\alpha\geq 0$, $\alpha x\in C$
Convex set A set $S$ such that for any $x,y\in S$ and $\alpha\in[0,1]$, $\alpha x+(1-\alpha)y\in S$
Convex hull The smallest convex set containing a given set of points
Graph Embedding Residual Vector A vector representing the additional error in graph fitting due to non-linear components. For details, refer to the related 5
@@
Symbol Description
$G_{ori}$ The original graph without prompting
$P_{\omega}$ Graph prompting method with parameter $\omega$
$C(G_{ori})$ The optimal embedding vector for the downstream task. $C(\cdot)$ can be understood as the optimal downstream task model
$G_{bri}$ The bridge graph that can be used to obtain $C(G_{ori})$ using the original model
$G_{p}$ The graph after prompting
$G_{\omega}$ The graph after prompting with parameter $\omega$
$\epsilon$ Represents the extent of Bridge set expansion, i.e. the “error”
$n$ Generally represents the number of layers in the GNN
$F$ Represents the dimension of the graph feature vector
$N$ Represents the number of nodes in the graph
$M$ Represents the number of graphs in the dataset $\Omega$
$\Phi$ Represents the aperture of the convex cone can be understood as the maximum opening angle
@@ A.2 Fundamental Lemmas
Here and in subsequent related content, we use $F_{\theta}$ to represent a GNN model whose aggregation process can be described by a diffusion matrix $S$. Here, $S$ is derived from the model type and the adjacency matrix $A$. For instance, in GCN, $S=A+\epsilon I$. The graph aggregation process is described as:
@@
$F_{\theta}(G_{\omega})=\text{ReadOut}(\sigma(S_{\omega}(\cdots\sigma(S_{\omega}X_{\omega}W)\cdots)W))=\text{ReadOut}(H^{(n)})$ (20)
@@ A.2.1 On the Range of Graph Embedding Matrix After Graph Prompting, one prompt node case
$S_{\omega}=S,\quad X_{\omega}=X+\mathbf{1}_{N}\mathbf{p}^{\top}$ (22)
Where $N$ is the number of nodes in the graph, $p$ is the prompt vector, $p\in\mathbb{R}^{F}$, $F$ represents the dimension of the parameter vector, $\mathbf{1}_{N}$ is a vector $\in\mathbb{R}^{N}$ with every component to be $1$.
@@
$S_{\omega}=\begin{pmatrix}S&l\\ l^{\top}&S_{NN}\end{pmatrix},\quad X_{\omega}=X+\mathbf{e}_{N}\mathbf{p}^{\top}$ (23)
Where $l\in\mathbb{R}^{N-1}$ is a column vector, $l_{i}\geq 0$, $\forall i\in\{1,\cdots,N-1\}$. In summary, the transformation of $X_{\omega}$ can be denoted as:
$X_{\omega}=X+\mathbf{c}\mathbf{p}^{\top},\quad\text{where }\mathbf{c}_{i}\geq 0$ (24)
Where $\mathbb{c}$ can be referred to as the coefficient vector.
@@
$\sigma(R_{i}^{\prime\top}+\Delta R+c_{i}\mathbf{p}^{\top})=\sigma(R_{i}^{\prime\top}+c_{i}\mathbf{p}^{\top}+c_{i}\mathbf{p}_{0}^{\top})$ $=\sigma(c_{i}(\mathbf{p}^{\top}+\mathbf{p_{0}}^{\top}))$
@@
$=\sigma(S_{\omega}(X+\mathbf{c}\mathbf{p}^{\top})W)$
$=\sigma(S_{\omega}XW+S_{\omega}\mathbf{c}\mathbf{p}^{\top}W)$
@@
Where $R=S_{\omega}XW$, each row of $R$ is a vector in the row space $R(W)$. $\mathbf{c}^{\prime}=S_{\omega}\mathbf{c}$, $c^{\prime}_{i}\geq 0$, $\mathbf{p}^{\top}W=\mathbf{p}^{\prime\top}$. Here, $\mathbf{p}^{\prime}$ spans $R(W)$ since $\mathbf{p}$ spans $\mathbb{R}^{N}$. ($c^{\prime}_{i}\geq 0$ since each element of $S_{\omega}$ is non-negative, the sum of each column in $S_{\omega}$ is strictly greater than 0 and the initial value of $\mathbf{c}$ is either $\mathbf{e}_{N}$ or $\mathbf{1}_{N}$. Note that the number of positive terms increases in $\mathbf{c}$.)
@@ A.2.2 On the Range of Graph Embedding Matrix After Graph Prompting, multiple prompt node case
$=\sigma(S_{i,\omega}(X_{i,\omega}+\mathbf{1}_{N}Q_{i}^{\top}P)W)$
@@
$\sigma(S_{i,\omega}X_{i,\omega}W)=\sigma(S_{i,\omega}(X_{i}+\mathbf{1}_{N}Q_{i}P)W)=\hat{R}^{\prime}+\mathbf{c}_{1,i}p^{\top}$
@@
$K_{\theta}(G_{i,\omega})=H_{i}^{(n)}=\hat{R}^{(n)}+\mathbf{c}_{n,i}p_{i}^{\top}=\hat{R}^{(n)}+\mathbf{c}_{n,i}Q_{i}^{\top}P$

### v2

@@ 5.3 On mapping to $\epsilon\text{-}B_{G}$ with single graph
(a) Error w.r.t matrices rank.
(b) Feature Number vs. $\epsilon$
(c) Graph Size vs. $\epsilon$
(d) Layer Number vs. $\epsilon$
@@
The parameters of interest include weight rank (Figure 3a), node feature dimension (Figure 3b), graph size (Figure 3c), and model layer number (Figure 3d). Here All-in-One-Plus extends All-in-One model by treating the inserting pattern as free parameters. and GPF-plus extends GPF by adding more than one tokens. Intuitively, with the increase of data complexity (e.g., larger features and graph size), the upper bound becomes larger in general. As the graph model becomes more complicated (e.g., layer number increase), the projected space becomes larger making the error bound intend to be smaller. When weight rank declines, the model’s capacity intends to poor results, making the error bound increase. More advanced graph prompts (e.g., All-in-One) usually have a lower error bound than the naive one (e.g., GPF). These observations can be naturally inferred from our theoretical analysis in section 4.1.
@@ 5.4 On mapping to $\epsilon\text{-}B_{G}$ with multiple graphs
Theorem 6 discusses a lower bound on the RMSE over the errors on multiple graphs with a single prompt token. In this section, we conducted experiments on the number of graphs in the dataset w.r.t the empirical minimum error. As shown in Figure 4, the minimum error shows an upward trend and then tends to saturate, which is highly consistent with the findings in Theorem 6.
Figure 4: $\epsilon$ range with a simple prompt token
Theorem 7 suggests that for the graph prompt with multiple tokens and multiple graphs, a small $k$ tokens is sufficient to achieve good performance. In particular, we wish to see how the error (loss) changes as $M$ (number of graphs) increases while $k$ remains fixed, and its counterpart case: how does the error change as $k$ increases while $M$ remains fixed? Here we explored the relationship between the number of prompt tokens, the number of graphs in the dataset, and the error. We present experimental results in Figure 5a and Figure 5b, which indicate two surfaces. From these figures we can find that both GPF and All-in-One show similar effects: when the number of prompt tokens exceeds $10$, the error becomes relatively small. As the number of prompt tokens increases further, the loss does not significantly decrease. Similarly, when the number of graphs increases and the number of prompt tokens is large, the decrease in error is also not that obvious.
(a) GPF Plus
(b) All in one
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgments
This research is supported by project #MMT-p2-23 of the Shun Hing Institute of Advanced Engineering, The Chinese University of Hong Kong, and by grants from the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK 14217622).
@@ A.1.1 Preliminaries
Graph Convolutional Networks (GCN) (Kipf & Welling, 2016) perform convolution operations on graph-structured data by aggregating feature information from a node’s neighbors. The recursive update rule for a GCN layer is:
$\mathbf{H}^{(i+1)}=\sigma\left(\tilde{\mathbf{A}}\mathbf{H}^{(i)}\mathbf{W}^{(i)}\right),$ (6)
where $\tilde{\mathbf{A}}=\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ is the normalized adjacency matrix, $\mathbf{H}^{(i)}$ is the node feature matrix at layer $i$, $\mathbf{W}^{(i)}$ is the learnable weight matrix, and $\sigma$ is a nonlinear activation function.
Graph Attention Networks (GAT) (Veličković et al., 2017) enhance GCNs by introducing attention mechanisms to weigh the importance of neighboring nodes during aggregation. In the simplest form of self-attention, the attention coefficient between node $j$ and node $k$ is based on the inner product of their feature vectors:
$e_{jk}$ $=\mathbf{H}_{j}^{(i)\top}\mathbf{H}^{(i)}_{k},$ (7)
$\alpha_{jk}$ $=\frac{\exp(e_{jk})}{\sum_{m\in\mathcal{N}(j)}\exp(e_{jm})},$ (8)
$\mathbf{H}_{j}^{(i+1)}$ $=\sigma\left(\sum_{k\in\mathcal{N}(j)}\alpha_{jk}\mathbf{H}_{k}^{(i)}\mathbf{W}^{(i)}\right),$ (9)
where $\mathcal{N}(j)$ denotes the neighbors of node $j$, $\alpha_{jk}$ is the normalized attention coefficient, and $\mathbf{W}^{(i)}$ is the learnable weight matrix at layer $i$, $\mathbf{H_{j}}$ is the embedding vector of the $j$-th node. This mechanism allows the model to focus on the most relevant neighbors when updating node representations.
@@
Pre-training and Fine-tuning (Zhu et al., 2021)involves two stages: first, a model is pre-trained on a large dataset to learn general representations, and then it is fine-tuned on a specific downstream task.
@@
$F_{\theta^{*}\rightarrow\theta^{\#}}(G_{\text{ori}})\to C(G_{\text{ori}}).$ (10)
@@
Graph Prompt Frameworks introduce learnable modifications to input graphs to enhance the performance of frozen pre-trained GNNs on downstream tasks. Two primary frameworks are prompt token vectors like GPF/GPF-Plus (Fang et al., 2022) and prompt subgraph like All-in-One (Sun et al., 2023a).
GPF (Graph Prompt Feature) adds a prompt vector to each node’s feature vectors. Let $\boldsymbol{\omega}=\{p\}$, $p\in\mathbb{R}^{F\times 1}$ be the learnable prompt vector, then the updated node features are:
@@
$p^{*}=\arg\min_{p}\sum_{G\in\mathcal{G}}\mathcal{L}_{T_{\text{dow}}}\left(F_{\theta^{*}}(P_{\boldsymbol{\omega}}(G))\right)$ (13)
GPF-Plus adds a combination of multiple prompt vectors to each node’s features, with different combinations for different graphs.. Let $\boldsymbol{\omega}={p_{1},\cdots,p_{k},Q}$, where $p_{i}\in\mathbb{R}^{F\times 1}$ represents a learnable prompt vector, and $Q\in\mathbb{R}^{M\times k}$ is a matrix in which the $i$-th row corresponds to the combination coefficients of the $k$ prompt vectors for graph $i$. Here, $M$ denotes the number of graphs in the dataset $\Omega$. Let $P=\begin{pmatrix}p_{1}^{\top}\\ \vdots\\ p_{k}^{\top}\end{pmatrix}$. The node features are updated in such a way:
@@
All-in-One optimizes the prompt tokens and their connections to adapt the pre-trained model to downstream tasks without altering the model parameters. For computational convenience, in the proof section below, we often use the equivalent form of All-In-One similar to Equation:23.
@@
Diffusion Matrix (Gasteiger et al., 2019) plays a crucial role in representing the diffusion process on a graph. Specifically, many GNN architectures can be expressed using the following formulation:
@@ A.1.2 Definitions
{xtabular}
p0.3p0.65 Term Explanation
@@
Adjacency matrix A square matrix used to represent a finite graph, where $A_{ij}=1$ if there is an edge from vertex $i$ to vertex $j$, and 0 otherwise.
Diffusion matrix The matrix equivalent to graph aggregation in GNNs.
Span For a set $V$, we say “p spans $V$” if $p$ can take any value in $V$.
$i$-th order embedding matrix The embedding matrix after $i$ iterations of message passing and aggregation in a GNN.
Cone A set $C$ such that for any $x\in C$ and $\alpha\geq 0$, $\alpha x\in C$.
Convex set A set $S$ such that for any $x,y\in S$ and $\alpha\in[0,1]$, $\alpha x+(1-\alpha)y\in S$.
Convex hull The smallest convex set containing a given set of points.
Graph Embedding Residual Vector A vector representing the additional error in graph fitting due to non-linear components. For details, refer to the related 5.
Table 2: Glossary
{xtabular}
p0.1p0.85 Symbol Description
$G_{ori}$ The original graph without prompting.
$P_{\omega}$ Graph prompting method with parameter $\omega$.
$C(G_{ori})$ The optimal embedding vector for the downstream task. $C(\cdot)$ can be understood as the optimal downstream task model.
$G_{bri}$ The bridge graph that can be used to obtain $C(G_{ori})$ using the original model.
$G_{p}$ The graph after prompting.
$G_{\omega}$ The graph after prompting with parameter $\omega$.
$\epsilon$ Represents the extent of Bridge set expansion, i.e., the “error.”
$n$ Generally represents the number of layers in the GNN.
$F$ Represents the dimension of the graph feature vector.
$N$ Represents the number of nodes in the graph.
$M$ Represents the number of graphs in the dataset $\Omega$.
$\Phi$ Represents the aperture of the convex cone, which can be understood as the maximum opening angle.
@@ A.2 Fundamental Lemmas
Here and in subsequent related content, we use $F_{\theta}$ to represent a GNN model whose aggregation process can be described by a diffusion matrix $S$. Here, $S$ is derived from the model type and the adjacency matrix $A$. For instance, in GCN, $S=A+\tau I$, where $\tau$ controls the strength of the self-loop for gathering information from the node itself. The graph aggregation process is described as:
@@
$F_{\theta}(G_{\omega})$ $=\text{ReadOut}(\sigma(S_{\omega}(\cdots\sigma(S_{\omega}X_{\omega}W)\cdots)W))$ (19)
$=\text{ReadOut}(H^{(n)})$ (20)
@@ A.2.1 On the Range of Graph Embedding Matrix After Graph Prompting, one prompt node case
$S_{\omega}=S,\quad X_{\omega}=X+\mathbb{1}_{N}\mathbf{p}^{\top}$ (22)
Where $N$ is the number of nodes in the graph, $p$ is the prompt vector, $p\in\mathbb{R}^{F}$, $F$ represents the dimension of the parameter vector, $\mathbb{1}_{N}$ is a vector $\in\mathbb{R}^{N}$ with every component to be $1$.
@@
$S_{\omega}=\begin{pmatrix}S&l\\ l^{\top}&S_{(N+1)(N+1)}\end{pmatrix},\quad X_{\omega}=\begin{bmatrix}X\\ \mathbb{0}_{F}\end{bmatrix}+\hat{\mathbf{e}}_{N+1}\mathbf{p}^{\top}$ (23)
Where $l\in\mathbb{R}^{N}$ is a column vector, $l_{i}\geq 0$, $\forall i\in\{1,\cdots,N\}$, $\hat{\mathbf{e}}_{N+1}$ is a vector $\in\mathbb{R}^{N+1}$ with (N+1)-th component to be $1$ and others to be 00. $\mathbb{0}_{F}\in\mathbb{R}^{F}$ with every component to be 00. Please note that we can shift the number of nodes($N$): the original number of nodes at this point can be denoted as $N-1$, to maintain consistency format with GPF case.
In summary, the transformation of $X_{\omega}$ can be denoted as:
$X_{\omega}=\tilde{X}+\mathbf{c}\mathbf{p}^{\top},\quad\text{where }\mathbf{c}_{i}\geq 0$ (24)
Where $\mathbf{c}$ can be referred to as the coefficient vector, $\tilde{X}$ can be either $X$ or the natural extension of $X$: $\begin{bmatrix}X\\ \mathbb{0}_{F}\end{bmatrix}$.
@@
$\sigma(R_{i}^{\prime\top}+\Delta R+c_{i}\mathbf{p}^{\top})$ $=\sigma(R_{i}^{\prime\top}+c_{i}\mathbf{p}^{\top}+c_{i}\mathbf{p}_{0}^{\top})$
$=\sigma(c_{i}(\mathbf{p}^{\top}+\mathbf{p_{0}}^{\top}))$
@@
$=\sigma(S_{\omega}(\tilde{X}+\mathbf{c}\mathbf{p}^{\top})W)$
$=\sigma(S_{\omega}\tilde{X}W+S_{\omega}\mathbf{c}\mathbf{p}^{\top}W)$
@@
Where $R=S_{\omega}\tilde{X}W$, each row of $R$ is a vector in the row space $R(W)$. $\mathbf{c}^{\prime}=S_{\omega}\mathbf{c}$, $c^{\prime}_{i}\geq 0$, $\mathbf{p}^{\top}W=\mathbf{p}^{\prime\top}$. Here, $\mathbf{p}^{\prime}$ spans $R(W)$ since $\mathbf{p}$ spans $\mathbb{R}^{N}$. ($c^{\prime}_{i}\geq 0$ since each element of $S_{\omega}$ is non-negative, the sum of each column in $S_{\omega}$ is strictly greater than 0 and the initial value of $\mathbf{c}$ is either $\hat{\mathbf{e}}_{N}$ or $\mathbb{1}_{N}$. Note that the number of positive terms increases in $\mathbf{c}$.)
@@ A.2.2 On the Range of Graph Embedding Matrix After Graph Prompting, multiple prompt node case
$=\sigma(S_{i,\omega}(X_{i,\omega}+\mathbb{1}_{N}Q_{i}^{\top}P)W)$
@@
$\sigma(S_{i,\omega}X_{i,\omega}W)=\sigma(S_{i,\omega}(X_{i}+\mathbb{1}_{N}Q_{i}P)W)=\hat{R}^{\prime}+\mathbf{c}_{1,i}p^{\top}$
@@
$K_{\theta}(G_{i,\omega})$ $=H_{i}^{(n)}$
$=\hat{R}^{(n)}+\mathbf{c}_{n,i}p_{i}^{\top}$ $=\hat{R}^{(n)}+\mathbf{c}_{n,i}Q_{i}^{\top}P$